OpenStudy (anonymous):

Pineapples cost $2.50 each and mangoes cost $1.50 each. Tim purchased 15 pieces of fruit for $29.50 How many pineapples and how many mangoes did he purchase?.

OpenStudy (anonymous):

x + y = 15 2.50x+1.50y = 29.50

OpenStudy (valpey):

My favorite approach is to use Hally85's equation and multiply the entire first line by -1.50. Then, when I add the two equations together, the y's cancel out.

OpenStudy (anonymous):

so multiply 29.50 by -1.50?

OpenStudy (anonymous):

No multiply that number to the first equation

OpenStudy (anonymous):

This is so that the y's are able to cancel each other out since it will be -1.5y and +1.5y

OpenStudy (anonymous):

that gave me 22.50, Im so sorry im no good at word problems.

OpenStudy (anonymous):

That's alright, you're doing it correctly. Just take it slowly. Ok so what do you get when you distribute -1.50(x+y=15)

OpenStudy (anonymous):

What you do to one part of the equation, you have to do to the other parts, correct? And that's why you distribute.

OpenStudy (anonymous):

Im still stuck????

OpenStudy (anonymous):

Okay well show me what you've got so far

OpenStudy (anonymous):

x+y=15

OpenStudy (anonymous):

Ok have you learned elimination and substitution?